Common radio of the following GP and the number of terms 10/7,1.…...343/1000
Answer by stanbon(75887)

10/7,1.…...343/1000
------
common ratio = r = 1/(10/7) = 7/10
-------
a(n) = a(1)*r^(n-1)
----
Solve for "n"::
343/1000 = (10/7)*(7/10)^(n-1)
---------
343/1000 = 1*(7/10)^(n-2)
----
(n-2)log(7/10) = log(343/1000)
----
n-2 = log(0.343)/log(0.7)
-------
n-2 = 3
Ans: n = 5
---------------